### v2

# Revisiting GW150914 with a non-planar, eccentric waveform model

Abstract
The first direct detection of gravitational waves by the LIGO collaboration, GW150914, marked the start of a new exciting era in astronomy, enabling the study of the Universe through a new messenger. Since then, the field has grown rapidly, with the development of increasingly more sophisticated techniques to detect, analyze and interpret the signals. In this paper we revisit GW150914, presenting updated estimates of its source parameters using a waveform model developed within the EOB formalism, able to describe gravitational-wave emission from generic non-circular, non-planar binaries. We provide a comprehensive analysis of the signal and its properties, considering and contrasting various scenarios for the source: from the simplest, aligned-spin quasi-circular binary black hole merger, to more complex scenarios, including precession, eccentricity or both. Unsurprisingly, we find that the signal is consistent with a quasi-circular ($e<0.08$ at $15$ Hz), slowly spinning $(\chi_{\rm eff}=-0.03^{+0.12}_{-0.13})$ binary black hole merger, a-posteriori validating a considerable body of works. This is the first analysis performed with an inspiral-merger-ringdown model containing both eccentricity and precession.
: Class. Quantum Grav.
1 Introduction
On September 14 2015 the LIGO-Virgo Collaboration (LVC[1, 2, 3] directly detected the first gravitational-wave (GW) signal from two compact objects, GW150914 [4]. To date, this event represents one of the loudest, best characterized GW signals. Since its discovery, GW150914 has been the subject of numerous studies across multiple domains of gravitational wave astronomy. Initial efforts were directed at characterizing the signal and its astrophysical properties [5, 6], using it to test the predictions of general relativity [7] and studying the properties of the binary black hole (BBH) population [8, 9]. These early analyses confirmed the consistency of GW150914 with general relativity (GR), identifying the signal as the coalescence of two approximately equal-mass, slowly spinning black holes merging along quasi-circular orbits into a final Kerr BH remnant with a mass of approximately $62M_{\odot}$ [6]. A multitude of subsequent studies have utilized GW150914 as a testbed for developing new techniques to analyze GW signals. These include – but are not limited to – advancements in parameter estimation [10, 11, 12, 13, 14], ringdown analyses [15, 16, 17, 18, 19, 20, 21, 22, 23], new tests of GR [24, 25, 26, 27, 28, 29], and improvements to waveform models [30, 31, 32, 33, 34, 35, 36]. Despite the application of increasingly refined techniques, the fundamental interpretation of the signal has remained largely unchanged.
In this paper, we revisit GW150914 presenting a comprehensive analysis of its properties. We employ TEOBResumS-Dalí, a physically complete model based on the Effective-One-Body (EOB) framework [37, 38], to study the signal considering a variety of scenarios for the source: from the simple coalescence of a planar, quasi-circular BBH system, to more complex scenarios including precession, eccentricity or, for the first time, the combination of the two.
The paper is structured as follows. In Sec. 2 we summarize the main effects that spin-induced precession and eccentricity have on a GW signal from BBH systems, describing the techniques used to model them and their implementation within TEOBResumS-Dalí. We then present a series of comparisons against selected numerical relativity (NR) simulations. In Sec. 3 we introduce the Bayesian parameter estimation (PE) framework used to analyze GW150914, and describe the different hypotheses considered in our analysis. Section 4 collects the main results of our analysis, discussing the posterior distributions of the parameters of interest and the implications for the astrophysical interpretation of the signal. Finally, we summarize our findings in Sec. 5.
Conventions
We denote the masses of the two BHs as $m_{1}$ and $m_{2}$, their mass ratio as $q=m_{1}/m_{2}\geq 1$ and their symmetric mass ratio as $\nu=q/(1+q)^{2}$. The total mass is $M=m_{1}+m_{2}$, while the chirp mass is $\mathcal{M}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}=\nu^{3/5}M$. We denote the dimensionless spin vectors of the two bodies as $\bm{\chi}_{i}=\bm{S}_{i}/m_{i}^{2}$, where $\bm{S}_{i}$ is the spin angular momentum of the $i$-th BH. The total angular momentum is then $\bm{J}=\bm{S}_{1}+\bm{S}_{2}+\bm{L}$, with $\bm{L}$ being the orbital angular momentum of the system. We define the effective spin as $\chi_{\rm eff}=(m_{1}\bm{\chi}_{1}+m_{2}\bm{\chi}_{2})\cdot\hat{\bm{L}}/M$, where $\hat{\bm{L}}$ is the unit vector along the direction of the orbital angular momentum, and the perpendicular spin parameter [39] as:
$\chi_{\rm p}=\max\left\{|\bm{\chi}_{1}^{\perp}|,\frac{4+3q}{4q^{2}+3q}|\bm{\chi}_{2}^{\perp}|\right\}\,$, (1)
where $|\bm{\chi}_{i}^{\perp}|$ are the projections of the spin vectors onto the plane perpendicular to $\bm{L}$. The waveform polarizations $_{+}, h×h_{\times}$ are decomposed in terms of the s=2s=-2 spin-weighted spherical harmonics $_{-2}Y_{lm}$ as
$h_{+}-ih_{\times}=\sum_{lm}h_{lm}{}_{-2}Y_{lm}(\iota,\phi_{\rm ref}),$ (2)
where $\iota$ and $\phi_{\rm ref}$ are the angles that define the orientation of the line of sight with respect to the inertial reference system of the binary at the frequency $f_{\rm ref}$. Unless otherwise specified, we use geometrized, mass-rescaled units in which $G=c=1$ and times are measured in multiples of the total mass $M$.
2 Modeling GWs from BBHs on generic orbits
The accurate modeling of GWs emitted by coalescing BBH systems represents a significant challenge due to the complex interplay of relativistic effects. Among these, spin-induced precession and orbital eccentricity play critical roles in shaping the phenomenology of a GW signal: capturing their influence requires sophisticated techniques.
Spin precession occurs when the spin vectors of the binary components are misaligned with the orbital angular momentum, inducing the precession of the individual spin vectors $\bm{S}_{1,2}$ and of the orbital angular momentum vector $\bm{L}$ around the total angular momentum $\bm{J}$ [40, 41, 42]. Consequently, the orbital plane itself oscillates, introducing characteristic modulations in the GW signal. The dominant emission occurs in a direction which is approximately that of the Newtonian orbital angular momentum [43, 44, 45, 46], which defines a special non-inertial “co-precessing” frame where the waveform closely resembles that of an aligned-spin system [47, 43, 44]. State of the art GW models track the evolution of this co-precessing frame via three Euler angles, and apply a time-dependent rotation (often referred to as the “twist”) to the co-precessing modes to reconstruct the waveform in a generic inertial reference frame.
Eccentricity (or, more generally, “non-circularity”), on the other hand, induces oscillations in the frequency and amplitude of the GW on the timescale of the orbit, with each pericenter passage accompanied by a “burst” of GW emission. Additionally, the presence of eccentricity reveals the well-known effect of periastron advance, which manifests in the waveform as modulations in the amplitude oscillations (see e.g. Fig. 4 of Ref. [48]). From a modeling perspective, the treatment of non-circular dynamics requires a generalization of the prescription used to describe GW emission and its dissipative effects. Historically, post-Newtonian (PN) expressions for the energy and momentum fluxes have been specialized to quasi-circular (QC) orbits, as eccentricity is efficiently radiated away [49], and thus was expected to be negligible once the system enters the LIGO-Virgo-Kagra Collaboration (LVK) frequency band. Energy and momentum fluxes for non-circularized binaries have been derived in PN theory up to 3 PN order (e.g. [50, 51, 52, 53]), and various examples of their factorizations have been tested in the point-particle limit and against NR simulations [54, 55, 56].
In this work, we use TEOBResumS-Dalí [37], a state of the art EOB model that combines PN and NR information to describe coalescing BBH systems. The formalism it is based on, EOB [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68], is a resummation of the PN dynamics of the system, including the effects of spins, precession, eccentricity, subdominant modes and tidal effects. The conservative Hamiltonian of TEOBResumS-Dalí incorporates point mass information up to 5PN [69, 70]. Spin-orbit contributions are included at next-to-next-to-leading order (NNLO[71, 70], and inverse-Taylor-resummed following the EOB prescriptions detailed in Ref. [72]. Even-in-spin effects are accounted for up to next-to-leading order (NLO). The radiative sector contains point-mass and spin terms according to Tab. I of Ref. [73], which also lists their resummation choices.
Within TEOBResumS-Dalí, we indirectly define the initial (model) eccentricity $e$ and anomaly $u$ as:
$r=\frac{p}{1+e\cos(u)}\,$, (3)
that is, the initial separation $r$ is given in terms of the semi-latus rectum $p$ (which can be determined from the initial average orbital frequency), the eccentricity $e$ and true anomaly $u$, related to the mean anomaly $\zeta$ via the eccentric anomaly and Kepler’s equation [74]. Non-circular effects are then accounted for via the prescription introduced in Ref. [75]: the quasi-circular leading order (LO) terms, both in the radiation reaction and waveform, are replaced with their exact analytical expressions valid on general orbits. Note that this simple technique has been validated against a large number of test-mass evolutions, proving to be more reliable with respect to the direct inclusion of PN expanded high order terms. Additionally, the radial radiation reaction force, usually set to zero for quasi-circular binaries, is non-negligible for binaries coalescing on generic orbits. It is included in the model as a resummation of the 2PN expression derived in Ref. [76]. Precession, on the other hand, is modeled via a hybrid PN-EOB scheme. We use the EOB orbital frequency $\Omega$ to drive the evolution of the orbit-averaged PN spin evolution equations of Ref. [77] up to merger. Beyond this point, identified as the last peak of the “pure” orbital EOB frequency (i.e., the term in $\Omega$’s expression coming from the orbital part of the EOB Hamiltonian, excluding the spin-orbit contribution), we implement a quasi-normal mode (QNM)-inspired prolongation of the Euler angles that describes the precession of the remnant BH’s emission [78, 32]. For more details, we refer the reader to Sec. IID of Ref. [32].
To both visually illustrate the impact of the various physical effects on the GW emission, as well as to display the faithfulness of our model, we now present three comparisons of TEOBResumS-Dalí with NR simulations: (i) SXS:BBH:0305 [79], which is a GW150914-targeted simulation with mass ratio $q=1.22$ and spins $\chi_{1}=0.33$, $\chi_{2}=-0.44$ aligned with the orbital angular momentum (Fig. 1); (ii) SXS:BBH:1389 [80], which is a spin-precessing simulation spanning more than 140 orbits having $q=1.63$, $\bm{\chi}_{1}=(-0.29,0.2,-0.3)$, $\bm{\chi}_{2}=(-0.1,0.42,0.16)$ (Fig. 2); (iii) RIT:BBH:1632 [81, 82], which is an eccentric, precessing BBH simulation with $q=1$ and $\bm{\chi}_{1}=\bm{\chi}_{2}=(0.7,0,0)$ (Fig. 3). For all of the simulations considered, we show the $(2,2)$ and $(2,1)$ modes, align the waveforms in the early inspiral and compute phase and amplitude differences all through to merger and ringdown. As expected, the quantitative agreement of our model with NR degrades as the complexity of the system increases. The phase difference at merger $\Delta\phi_{22}^{\rm EOBNR}=\phi_{22}^{\rm EOB}-\phi_{22}^{\rm NR}$, which amounts to $\sim-0.3$ rad for the quasi-circular, aligned spins system, grows to more than $1$ rad for the spin-precessing and eccentric, precessing cases. Nonetheless, qualitatively, the main features and oscillations induced by eccentricity and/or precession are fully captured by the model: from inspecting the $(2,1)$ mode from Fig. 2, we immediately observe that the four precession cycles that the system undergoes (as inferred from the number of minima of the amplitude of this mode) are reproduced by the simple prescription employed. Similarly, periastron passages (corresponding to the amplitude peaks of Fig. 3) are correctly captured as well.
A quantitative and comprehensive validation of our model against NR simulations has been performed in Ref. [38]. There, TEOBResumS-Dalí was compared with $1395$ NR simulations from the SXS [83], RIT [84, 85], CoRe [86, 87] and ICC catalogs [88], covering a wide range of masses, spins and orbital configurations. Anticipating the results of Sec. 3, we find that in the region of interest for GW150914 (that is, equal-mass, slowly spinning BBHs) our model is typically more than $99.9\%$ faithful to NR. The high eccentricity regime has been further tested in Ref. [89], where the model was shown to be in agreement with NR simulations with initial eccentricities as high as $e\sim 0.7$.
Figure 1: Comparison of the $(2,2)$ and $(2,1)$ modes of the TEOBResumS-Dalí model against the SXS:BBH:0305 simulation, which is a GW150914-targeted simulation with mass ratio $q=1.22$ and aligned spins $\chi_{1}=0.33$, $\chi_{2}=-0.44$. The (2,2) mode is aligned in the early inspiral, and the time and phase shifts obtained used also to align the $(2,1)$ mode. The small oscillations observed in the EOB/NR phase difference are due to a small residual eccentricity in the EOB waveform, related to the adiabatic initial conditions employed [34]. The mismatch [90] among the two waveforms is below $10^{-3}$ for masses compatible with GW150914.
Figure 2: Comparison of the $(2,2)$ and $(2,1)$ modes of the TEOBResumS-Dalí model against the SXS:BBH:1389 simulation, which is a spin-precessing simulation with mass ratio $q=1.63$, and spins $\bm{\chi}_{1}=(-0.29,0.2,-0.3)$, $\bm{\chi}_{2}=(-0.1,0.42,0.16)$ The $(2,2)$ mode is aligned in the early inspiral, and the time and phase shifts obtained used also to align the $(2,1)$ mode. Remarkably, the EOB and NR waveforms maintain phase coherence and a good amplitude agreement throughout the entire inspiral and merger, with the exception of the last few cycles, where the EOB model exhibits a phase lag of $\sim 1$ rad. The spin-induced precession of the system is clearly visible in the $(2,1)$ mode: nodes of the amplitude correspond to the instants at which the azimuthal angle between the inertial frame and the co-precessing frame is equal to 0. The mismatch [90] among the two waveforms is $\sim 5\times 10^{-3}$ for masses compatible with GW150914.
Figure 3: Comparison of the $(2,2)$ and $(2,1)$ modes of the TEOBResumS-Dalí model against the RIT:BBH:1632 simulation, which is an eccentric, precessing simulation with mass ratio $q=1$ and aligned spins $\chi_{1}=(0.7,0,0)$, $\chi_{2}=(0.7,0,0)$. Similarly to the case of SXS:BBH:1389, the $(2,2)$ mode is only weakly affected by the presence of precession. Eccentricity, on the other hand, induces very recognizable oscillations in the amplitude and frequency of the waveform associated to the radial motion (pericenter passages) of the binary. The mismatch [90] among the two waveforms is $\mathcal{O}(10{-2})$ for masses compatible with GW150914.
3 GW parameter estimation
Given some data $d$ collected by the detectors, the goal of GW PE is to estimate the probability distribution of the parameters $\bm{\theta}$ which, under some hypothesis $H$, describe a binary system given the data, $p(\bm{\theta}|d,H)$. This problem is typically tackled with the use of Bayesian inference techniques [91, 92, 12], according to which $p(\bm{\theta}|d,H)$ can be estimated via:
$p(\bm{\theta}|d,H)=\frac{p(d|\bm{\theta},H)p(\bm{\theta}|H)}{p(d|H)}.$ (4)
In the expression above, $p(d|\bm{\theta},H)$ is the likelihood function, which contains the information about the GW signal and the noise of the detectors; $p(\bm{\theta}|H)$ is the prior distribution, which encodes our assumptions about the distribution of the parameters before the data is observed, and is usually chosen to be uniform in the parameters of interest (“agnostic” prior), or motivated by physical models; $p(d|H)=\int d\bm{\theta}p(d|\bm{\theta},H)p(\bm{\theta}|H)$ is the evidence (or “marginal likelihood”), a constant that ensures the posterior is properly normalized. In the context of model selection, $p(d|H)$ is used to compute the Bayes’ factor (BF) between two models, which quantifies the relative probability of the data given the two models and is used to determine whether one is favored by the data:
$\log\mathcal{B}^{1}_{2}=\log p(d|H_{1})-\log p(d|H_{2})=\log\frac{p(d|H_{1})}{p(d|H_{2})}.$ (5)
We apply this framework to GW150914. We describe the system via a set of 17 parameters: the masses of the binary components $m_{1}$ and $m_{2}$, the (dimensionless) spin vectors of the components $\bm{\chi_{1}}$ and $\bm{\chi_{2}}$, the initial eccentricity of the orbit $e$ and the mean anomaly of the system $\zeta$, the luminosity distance of the source to the detectors $D_{L}$, the inclination of the orbit with respect to the line of sight $\iota$, the polarization angle $\psi$, the time of coalescence $t_{c}$ at which the signal reaches its peak amplitude [^1] More precisely, we use the last peak of the waveform invariant amplitude to define merger, see e.g. App. C of Ref. [93], the reference orbital phase $\phi_{\rm ref}$, and the sky position of the source $(\alpha,\delta)$. We sample in chirp mass $\mathcal{M}\in[20,40]M_{\odot}$ and (inverse) mass ratio $1/q\in[0.05,1]$, enforcing uniform priors on the component masses $m_{1},m_{2}$, constrained to be $\in[3,100]M_{\odot}$; over the whole sky area for the angles $(\alpha,\delta)$, impose uniform priors with periodic boundaries on $\psi\in[0,2\pi]$ and $\phi_{\rm ref}\in[0,2\pi]$, and employ a power-law prior for the luminosity distance $D_{L}\in[10,3000]$ Mpc. We use a flat prior for the time of coalescence $t_{c}\in[t_{0}-1.6\mathrm{s},t_{0}+1.6\mathrm{s}]$, where the reference GPS time is $t_{0}=1126259462.891$, but instead of sampling it we analytically marginalize over it when computing the likelihood. Rather than directly sampling the spin components, we decompose the spin vectors in terms of their magnitudes, their tilt angles with respect to the orbital angular momentum $\bm{L}$, the azimuthal angle separating the spin vectors $\phi_{12}$ and the azimuthal angle angle $\phi_{\rm JL}$. Further, rather than directly sampling in $\iota$, we sample the cosine of the angle $\theta_{\rm JN}$ between the line of sight and the total angular momentum $\bm{J}$ of the system. Given our aim of contrasting four different scenarios, each corresponding to a different orbital configuration, priors on $e,\zeta,\bm{\chi_{1}}$ and $\bm{\chi_{2}}$ depend on the chosen underlying hypothesis:
  • quasi-circular, aligned-spin case (“QC-AS”): we fix $e=0,\zeta=0$ and sample the spins using an “aligned spin” prior, considering magnitudes $|\bm{\chi}_{1}|\in[0,0.99]$, $|\bm{\chi}_{2}|\in[0,0.99]$. The full model is thus $10$-dimensional.
  • Quasi-circular, precessing-spin case (“QC-PS”): we fix $e=0,\zeta=0$ and use an isotropic spin prior. The full model is thus $14$-dimensional.
  • Non-circular, aligned-spin case (“NC-AS”): we use uniform priors on $e\in[0,0.4]$ [^2] These prior bounds are inspired by previous analyses of the event, which found it to be consistent with a low eccentricity merger [34, 94, 36] and $\zeta\in[0,2\pi]$, and sample over the spin parameters employing the same prior as the QC-AS case. The full model is thus $12$-dimensional.
  • Non-circular, precessing-spin case (“NC-PS”): we use uniform priors on $e\in[0,0.4]$ and $\zeta\in[0,2\pi]$, and an isotropic spin prior. The full model is thus $16$-dimensional.
We evaluate the posterior distribution $p(\bm{\theta}|d)$ using the bilby package [92], coupled with the nested sampling algorithm as implemented in dynesty [95]. We use $2048$ livepoints and $60$ as number of accepted steps (naccept), running the analysis on $48$ cores. We analyze $8$ seconds of data around the GPS time of the event, focusing on the frequencies between $20$ and $896$ Hz. We download the data from the Gravitational Wave Open Science Center (GWOSC) and use the publicly available power spectral density (PSD) and calibration envelopes as computed by the LVK collaboration [96]. We generate waveforms with TEOBResumS-Dalí using a sampling rate of $4096$ Hz, and consider the $(\ell,|m|)=(2,1),(2,2),(3,3)$ and $(4,4)$ co-precessing modes for the waveform generation. We use an initial orbit-averaged frequency for waveform generation of $15$ Hz, lower than the minimum one in the analysis. This choice is motivated by the fact that the instantaneous frequency of an eccentric waveform varies throughout each orbit. Therefore, if we were to generate waveforms starting at 2020Hz we would risk miss some signal power. As our model natively outputs waveforms in the time domain, each likelihood evaluation requires us to perform a Fast Fourier transform (FFT). With these settings, the analyses take between $9$ and $17$ days to complete, depending on the model complexity.
Parameter GWTC QC AS QC PS NC AS NC PS
$\mathcal{M}[M_{\odot}]$ $30.4^{+1.6}_{-1.7}$ $30.9^{+1.4}_{-1.4}$ $31.2^{+1.6}_{-1.6}$ $30.6^{+1.5}_{-1.5}$ $30.9^{+1.6}_{-1.6}$
$1/q$ $0.85^{+0.14}_{-0.22}$ $0.89^{+0.10}_{-0.17}$ $0.87^{+0.12}_{-0.22}$ $0.89^{+0.10}_{-0.17}$ $0.86^{+0.13}_{-0.21}$
$\chi_{\rm eff}$ $05^{+0.11}_{-0.15} 0.0$$3^{+0.10}_{-0.12} 0.0$$1^{+0.12}_{-0.13} 0.0$$6^{+0.11}_{-0.12} 0.0$$03^{+0.12}_{-0.13}$
$d_{L}$ [Mpc] $60^{+130}_{-140} 44$$40^{+160}_{-160}$ $480^{+140}_{-150}$ $430^{+160}_{-160}$ $470^{+130}_{-150}$
$\iota$ [rad] $2.63^{+0.36}_{-0.43}$ $2.61^{+0.38}_{-0.56}$ $2.79^{+0.26}_{-0.54}$ $2.61^{+0.39}_{-0.57}$ $2.80^{+0.26}_{-0.56}$
$e$ $^{+0.06}_{-0.04} 0.040.04$$^{+0.06}_{-0.04}$
$\zeta$ [rad] $3^{+2.7}_{-2.9} 3.3$$3^{+2.7}_{-3.0}$
$\chi_{\rm p}$ $0.51^{+0.38}_{-0.40}$ $0.41^{+0.41}_{-0.32}$ $0.40^{+0.42}_{-0.31}$
$\log{\mathcal{B}}^{S}_{N}$ $280.55^{+0.11}_{-0.11}$ $280.44^{+0.11}_{-0.11}$ $279.41^{+0.12}_{-0.12}$ $279.29^{+0.12}_{-0.12}$
Table 1: Summary of the parameters of GW150914 from GWTC-2.1 (first column) and under the four hypotheses considered in this work. The (natural) logarithm of the BF comparing the signal to the noise hypothesis $\log\mathcal{B}^{S}_{N}$ is reported in the last row, confirming the expectation that this event is most consistent with a quasi-circular inspiral of two BHs, and that non-circularity is disfavoured.
Figure 4: Marginalized two- and one- dimensional posterior distributions for the common parameters of GW150914 in the four scenarios considered. We additionally compare our results to those from the GWTC catalog paper [97], here shown as the black contours. We find broad agreement between all analyses, with GWTC-2.1 predicting slightly lower chirp mass and inverse mass ratio
4 Results
Fig. 4 shows the posterior distributions for the common intrinsic parameters of GW150914 in the four scenarios considered. Most notably, the masses and effective spin are broadly consistent among all analyses, with $\mathcal{M}=30.90^{+1.63}_{-1.59}$, $1/q=0.86^{+0.13}_{-0.21}$ and $\chi_{\rm eff}=-0.03^{+0.12}_{-0.13}$ for the non-circular, precessing-spin model. The values recovered are also in agreement with those reported by past studies performed with different models [30, 33, 35, 36], as well as from GWTC-2.1 [97]. The nominal model eccentricity (left panel of Fig. 5) is constrained for both the planar and precessing spins scenarios to $e=0.04^{+0.06}_{-0.04}$ at $90\%$ credibility at $15$ Hz, showing considerable support towards the left boundary $e=0$. The anomaly parameter, instead, cannot be measured and closely follows the flat prior employed, confirming the expectation of a quasi-circular binary. Finally, in-plane spin components (right panel of Fig. 5) are also not well constrained, with $\chi_{\rm p}=0.40^{+0.42}_{-0.31}$ spanning the width of the entire prior distribution. With respect to the distrubution reported in GWTC-2.1, we find a slight shift towards lower values of $\chi_{\rm p}$, which is however consistent with, e.g., the results of Ref. [35] and points towards small differences due to the modeling of spin precession.
When comparing the different hypotheses in terms of log Bayes factors, we do not find any scenario to be strongly favored over the others (see Tab. 1). However, we do find that adding non-circularity to the model tends to decrease its Bayesian evidence, with the QC-AS and QC-PS models being favored over their NC counterparts by $\log\mathcal{B}^{\rm QC}_{\rm NC}\sim 1$. As such, we conclude that the data is consistent with a quasi-circular BBH merger, with small effective spin and no strong evidence in favor of precession. To visually confirm our result, we show the time-domain reconstruction from the non-circular, precessing-spin model in Fig. 6, overlaid on the whitened strain data from the Hanford and Livingston detectors. Compared to the simulations considered in the previous section, the (whitened) reconstructed waveforms show no clear signatures of eccentricity or precession.
Figure 5: Marginalized two- and one- dimensional posterior distributions for the spin parameters $\chi_{\rm eff}$ and $\chi_{p}$ (left) and the eccentricity $e$ and mean anomaly $\zeta$ (right). While $\chi_{\rm eff}$ is well constrained and in all case consistent with zero, $\chi_{p}$ spans the entire prior. The eccentricity at $15$ Hz is well constrained, with $e<0.08$ at $90\%$ credibility. Finally, the mean anomaly is follows the uniform prior employed. This is consistent with the expectation of a quasi-circular binary, where no value of $\zeta$ is preferred.
Figure 6: Whitened strain data from the Hanford (top) and Livingston (bottom) detectors, with the prediction from the non-circular, precessing-spin model overlaid; the thick line corresponds to the maximum likelihood waveform, the colored region is the 90% equal-tailed credible interval. As can be seen, the signal is visually consistent with a QC BBH merger, with no clear evidence of eccentricity or precession. This is reflected in the fact that the bulk of the posterior distribution for this model has low effective and precessing spin parameters, and eccentricity consistent with zero.
5 Conclusions
Despite the fact that GW150914 has been the subject of extensive research since its discovery, its usefulness as a testbed for new techniques and models has not diminished throughout the years. Its re-analysis as presented in this work provides a significant step forward in the field of GW astrophysics. First, the findings presented a-posteriori validate a considerable body of past research, performed assuming that GW150914 is a quasi-circular BBH merger. Second, they represent the first instance of a full GW PE performed with an inspiral-to-ringdown non-planar, non-circular model. This achievement underscores the growing sophistication of GW modeling and PE techniques, highlighting the importance of incorporating more realistic, complex waveforms when interpreting observational data.
Nonetheless, more work is needed to further improve GW models in view of next generation space- and ground- based detectors. Beyond improvements in their accuracy, additional physical effects should be included in the waveform generation. These include higher order harmonics beyond those considered here, including $m=0$ modes which encode GW memory [98, 99, 100, 101, 102, 103, 104], improved descriptions of tidal heating and torquing [105, 106, 107], as well as phenomenological descriptions of environmental effects [108, 109, 110]. The latter, while not expected to be relevant for GW150914 [111], may be important for the analysis of future events, especially those originating from dense stellar environments [112, 113, 114, 115, 116, 117]
The increased computational cost of TEOBResumS-Dalí due to the inclusion of both eccentricity and precession is a clear limitation for the analysis of large catalogs of GW events. Various avenues can be explored to address this issue. On the waveform generation side, the use of surrogate models [118, 119, 80, 120, 121, 122, 123] can significantly speed up the generation of the waveforms. Preliminary results in this direction have been presented, for nonspinning [124, 125] and aligned-spin [126, 127] eccentric binaries. Recently, a surrogate model for aligned-spin TEOBResumS-Dalí was constructed [123], employing a further decomposition of each harmonic mode of emission into modes with monotonically increasing frequency. Alternatively, reduced order modeling techniques [118, 128, 129, 130, 122] or analytical approximations [131, 132, 133] can also be employed to speed up the generation of the waveforms. On the PE side, the use of more efficient sampling algorithms [134, 135], deep-learning techniques [11, 136] used as high-quality posterior proposals [13] or approximate or marginalized likelihoods [137, 138, 139] can significantly reduce the computational cost of the analysis; for a review, see Ref. [140].
The ability to perform PE on systems where orbital eccentricity and spin precession play significant roles opens the door to more detailed and comprehensive studies of the BBH population, as the two effects are closely linked to the formation channels of BHs in dynamical environments. For instance, BBHs formed in galactic fields are generally expected to exhibit spins aligned with the orbital angular momentum and to inspiral along quasi-circular orbits [141, 142, 143, 144, 145]. In contrast, BBHs originating from dense stellar environments are more likely to display significant spin misalignments and residual eccentricities due to frequent dynamical interactions [146, 147, 148, 149]. As gravitational-wave detectors continue to increase in sensitivity, future analyses utilizing waveform models such as TEOBResumS-Dalí will provide deeper insights into the properties of BBH systems, shedding light on their formation pathways and evolutionary histories.
Acknowledgments
The authors would like to thank S. Bernuzzi and A. Nagar for useful discussions. RG is grateful to K. Chandra and I. Gupta for constant support and encouragement throughout the development of this work, and to M. Poulsen, J. Larsen and K.B. Larsen for inspiration. The version of TEOBResumS-Dalí employed in this work is available at https://bitbucket.org/teobresums/teobresums/src/2505.21612/. RG acknowledges support from NSF Grant PHY-2020275 (Network for Neutrinos, Nuclear Astrophysics, and Symmetries (N3AS)). JL and DC acknowledge support from the Italian Ministry of University and Research (MUR) via the PRIN 2022ZHYFA2, GRavitational wavEform models for coalescing compAct binaries with eccenTricity (GREAT). ST acknowledges support from NSF grant NSF PHY-2207780. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The authors are grateful for computational resources provided by LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459.